# Calibrated Physics-Informed Uncertainty Quantification

### v1

@@ Abstract
Neural PDEs offer efficient alternatives to computationally expensive numerical PDE solvers for simulating complex physical systems. However, their lack of robust uncertainty quantification (UQ) limits deployment in critical applications. We introduce a model-agnostic, physics-informed conformal prediction (CP) framework that provides guaranteed uncertainty estimates without requiring labelled data. By utilising a physics-based approach, we are able to quantify and calibrate the model’s inconsistencies with the PDE rather than the uncertainty arising from the data. Our approach uses convolutional layers as finite-difference stencils and leverages physics residual errors as nonconformity scores, enabling data-free UQ with marginal and joint coverage guarantees across prediction domains for a range of complex PDEs. We further validate the efficacy of our method on neural PDE models for plasma modelling and shot design in fusion reactors.
@@ 1 Introduction
Numerical PDE solvers are essential tools in scientific and engineering simulations (Danabasoglu et al., 2020; Giudicelli et al., 2024), yet their computational demands and environmental impact remain significant challenges (Horwitz, 2024). Machine learning approaches have emerged as efficient alternatives (Bertone et al., 2019; Karniadakis et al., 2021), successfully deployed across weather forecasting (Lam et al., 2023; Kurth et al., 2023), fluid dynamics (Jiang et al., 2020; Pfaff et al., 2021), and nuclear fusion applications (Poels et al., 2023; Carey et al., 2024; Gopakumar & Samaddar, 2020). Neural PDE solvers provide rapid approximations but present a critical cost-accuracy trade-off. While generating outputs consistently, their solutions may violate physical constraints or produce misleading results with high confidence (Gopakumar et al., 2023a). A typical neural-PDE framework (see Figure 1) trains surrogate models on numerical solver data to predict PDE evolution under various conditions, with uncertainty quantification (UQ) methods reverting to numerical solvers when predictions fail coverage thresholds. However, current UQ methods lack statistical guarantees (Zou et al., 2024), require extensive simulation data (Gopakumar et al., 2024a), or demand architectural modifications (Abdar et al., 2021).
Figure 1: Neural PDE framework: Neural PDE solvers use data from traditional numerical solvers to quickly approximate PDEs across various conditions (shown by black arrows). To ensure reliability, these models incorporate uncertainty quantification (UQ). If the predicted error exceeds a coverage threshold $\epsilon$, the numerical solver is used and adds to training data; otherwise, predictions are used as output (shown by red arrows). This paper focuses on developing a new UQ method for assessing confidence in neural PDE models within the gray-shaded region.
To address these limitations, we propose a framework combining PDE residuals over neural PDEs with Conformal Prediction (CP) to provide uncertainty estimates that guarantee coverage. Our approach evaluates Physics Residual Errors (PRE) from neural PDE solver predictions and performs calibration using marginal and joint CP formulations. Our method provides statistically valid coverage within the residual space (Vovk et al., 2005), offering error bounds based on physical conservation law violations. The framework is model-agnostic and requires no additional data. It yields interpretable uncertainty bounds indicating the model’s physical inconsistencies, addressing the (over)confidence issue of neural PDE solvers (Zou et al., 2024). Our contributions:
@@ 2 Related Work
Recently, CP, as a method of performing UQ, has been gaining popularity for usage with spatio-temporal data Sun (2022). Several works have explored the inductive CP framework for spatial and sequential data (Stankeviciute et al., 2021; Xu & Xie, 2021; Xu et al., 2023), including in the operator space (Ma et al., 2024). In Gopakumar et al. (2024a), the marginal-CP framework is extended to pre-trained as well as to fine-tuned surrogate models for physical system modelling across an infinite-dimensional setting. Alternatively, error bounds for PDE surrogates have been devised by Gray et al. (2025) using set propagation to project the singular value decomposition of the prediction error to the prediction space.
@@
The usage of PDE residuals under the guise of Physics-Informed Machine Learning (PIML) (Karniadakis et al., 2021) was made popular as an optimisation strategy for Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019) and has found application in optimising neural operators (Li et al., 2024) and soft/hard enforcement of the physical constraints to deep learning models (Du et al., 2024; Chalapathi et al., 2024). However, they have rarely been used as a tool for providing UQ to the surrogate models, and where they have found application, UQ remained uncalibrated (Zhu et al., 2019). The majority of literature in UQ for neural PDE solvers has been looking at Bayesian methods, such as dropout, Bayesian neural networks, and Monte Carlo methods (Geneva & Zabaras, 2020; Zou et al., 2024; Psaros et al., 2023), which lack guarantees or are computationally expensive.
@@ 3.1 Neural PDE Solvers
Here, $X$ defines the spatial domain bounded by $\Omega$, $[0,T]$ the temporal domain, $D_{X}$ and $D_{t}$, the composite operators of the associated spatial and temporal derivatives. The PDE is further defined by the boundary condition $g$ and initial condition $a$, which can be parameterised by $\lambda$. The set of solutions of field variables are expressed as $u\in\mathcal{U}$.
Neural PDE solvers as surrogate models aim to learn the behaviour governed by Equation 1 using a parameterised neural network $\mathcal{NN}_{\theta}$. Starting from the initial conditions, the network is trained to solve the spatio-temporal evolution of the fields given by $\Omega\,\cup\,[0,T]$. Neural operators $\mathcal{NO}_{\theta}$ are a special class of neural networks that learn the operator mapping from the function space of the PDE initial conditions $a\in\mathcal{A}$ to the function space of solutions $u\in\mathcal{U}$. A neural operator for solving an initial-value problem can be expressed as
@@
A Fourier Neural Operator (FNO) is an autoregressive neural operator that learns the spatio-temporal evolution of PDE solution by leveraging the Fourier transform as the kernel integrator (Li et al., 2021). The field evolution is learned using tuneable weight matrices of the network, parameterised directly in the Fourier space of the PDE solutions.
Since CP and our extension of it provide a post-hoc measure of quantifying the uncertainty of a neural PDE, it remains agnostic to model choice and training conditions. Considering the model independence of our approach, we restrict our experiments to modelling PDEs with an FNO. The FNO is chosen due to its cost-accuracy trade-off and efficiency as demonstrated by de Hoop et al. (2022) and Gopakumar et al. (2023b). CP over a range of neural-PDE solvers has been applied by Gopakumar et al. (2024a), who also demonstrate that the coverage guarantees are upheld irrespective of the model choice, not needing us to experiment with various model architectures.
@@ 3.2 Conformal Prediction
Conventionally, nonconformity scores act on the model predictions and a labelled dataset (Kato et al., 2023). For deterministic models, they are often formulated as the Absolute Error Residual (AER) of the model predictions $\hat{f}(X)$ and targets $Y$. For probabilistic models, the score function (STD) is the absolute error of the prediction means $\hat{f_{\mu}}(X)$ and the targets $Y$, normalised by the standard deviation of the prediction $\big{(}\hat{f_{\sigma}}(X)\big{)}$. Having obtained a distribution of nonconformity scores $\hat{s}$ of the calibration dataset $(X_{i},Y_{i})_{i=1}^{n}$, a quantile $\hat{q}$ corresponding to the desired coverage $1-\alpha$ is estimated from its cumulative distribution function $F_{\hat{s}}$ (Papadopoulos, 2008):
@@
The quantile estimates the error bar associated with desired coverage and is combined with the new prediction to obtain the prediction sets. The nonconformity score functions and their prediction sets for AER and STD are given in Table 1.
@@ 4 Physics Residual Error (PRE)
We introduce a novel data-free nonconformity score based directly on the PDE for surrogate models. The Physics Residual Error (PRE) is defined as the PDE residual (Saad & Schultz, 1986) estimated over the discretised PDE solution obtained from the surrogate model. For an abstract PDE as in Equation 1, the PDE residual is the evaluation of the composite differential operator $D$. The PDE residual is treated as a score function by taking its L1 norm as indicated in Table 1. While well-defined PDEs have solutions obeying Equations 1, 2 and 3, numerical solutions often fail to converge to the true solution (Pinder, 2018). Neural PDEs, trained on approximate numerical data, are further prone to non-convergent predictions. In PDE numerical analysis, the norm of the PDE residual is often used as a criterion for stability, convergence, and accuracy (Iserles, 2009). The PRE typically represents the violation of conservation laws associated with the physical system. Using the residual error as a nonconformity score evaluates the neural PDE solver’s performance by quantitatively estimating its effectiveness in modelling a PDE, leading to guaranteed coverage bounds within the residual space.
@@
The norm $\left|D(\mathcal{NO}_{\theta}(u))\right|$ of the residual operator itself provides a measure of UQ for the neural PDE. However, it is limited by the accuracy of the gradient estimation method and can become computationally expensive when exploring a vast solution space (Tolsma & Barton, 1998). By using the residual norm as a nonconformity score, we further calibrate the approximate physics residual error that is obtained by an inexpensive and coarse differential operator. CP using PRE provides statistically valid and guaranteed error bars across the PDE’s residual space, incorporating physical information into the calibration procedure and providing a calibrated measure of the physical misalignment of the surrogate model.
PRE as a nonconformity score enables data-free conformal prediction. The estimated scores rely only on the neural PDE predictions over a range of initial conditions, not on the target as in AER and STD. The only criterion is that the calibration and prediction domains arise from exchangeable initial conditions. As shown in Table 1, PRE gives “prediction sets” independent of the prediction inputs (see Appendix B for formalism). Traditional CP methods rely on calibration using observed data to construct prediction intervals that contain the true values with a specified confidence level $\alpha$. These methods guarantee that the true solution will lie within the estimated error bounds based on this empirical calibration. In contrast, our PRE-CP formulation takes a fundamentally different approach. Instead of requiring target data for calibration, we leverage the unique property of PDEs where the true solution in the residual space can be treated as zero. This eliminates the need for empirical calibration data altogether. Our method focuses on ensuring that predictions themselves fall within coverage bounds $\mathbb{C}^{\alpha}$, rather than guaranteeing that the true solution lies within these bounds. This allows us to validate prediction sets without access to ground truth data - a significant advantage over traditional CP approaches. We formalize this novel property in our theoretical framework presented in Appendix A.
@@ 4.1 Marginal-CP
The CP formulation was initially conceptualised for calibrating univariate functions with single-point outputs (Vovk et al., 2005). It has recently been extended to spatio-temporal data, with multi-dimensional outputs with an immutable tensor structure (Gopakumar et al., 2024a). Within such spatio-temporal settings, CP has been implemented to provide marginal coverage, i.e. the calibration procedure provides independent error bars for each cell within the spatio-temporal domain. For an output tensor $Y\in\mathbb{R}^{N_{x}\times N_{y}\times N_{t}}$, where $N_{x},N_{y},N_{t}$ represent the spatio-temporal discretisation of the domain, marginal-CP uses the non-conformity scores outlined in Tables 1, 1 and 1 across each cell of $Y$ to obtain error bars, which will be compliant with Equation 5 for each cell. Marginal-CP using PRE helps indicate regions within a single prediction that lie outside the calibrated bounds of physics violation and require specific attention, treating those predictions with caution.
@@ 4.2 Joint-CP
The joint-CP formulation constructs a calibration procedure that provides coverage bands for multivariate functions. These coverage bands expand across the entire simulation domain $\Omega\times[0,T]$ (discretised as $\mathbb{R}^{N_{x}\times N_{y}\times N_{t}}$) rather than an individual cell within it. For a coverage band $\mathbb{C}^{\alpha}$, the joint-CP formulation ensures that $1-\alpha$ predictions/solutions lie within the bounds. For performing joint-CP, the non-conformity scores are modified to reflect the supremum of the score functions $\mathcal{S}$ in Tables 1, 1 and 1. They are modulated by the standard deviation $\sigma$ of the calibration scores (Diquigiovanni et al., 2021) to obtain prediction bands with varying widths based on local behaviour (Diquigiovanni et al., 2022). The modifications of the score functions and prediction sets to perform CP are given by
$\hat{s}$ $=\sup_{X\in\Omega,\,t\in[0,T]}\bigg{(}\frac{\mathcal{S}}{\sigma(\mathcal{S})}\bigg{)},$ (7)
$\mathbb{C}^{\alpha}$ $=\mathcal{P}\pm\;\hat{q}^{\alpha}\cdot\sigma(\mathcal{S}),$ (8)
where $\mathcal{S}$ and $\mathcal{P}$ are the formulations of the nonconformity scores and the prediction at $X_{n+1}$ used for marginal-CP as shown in Tables 1, 1 and 1. Joint-CP becomes particularly useful in identifying predictions that fail to fall within coverage, allowing us to accept or reject a prediction based on a predetermined probability. Similar to that demonstrated by Casella et al. (2004), our framework can perform acceptance-rejection using a CP-based criterion. The acceptance probability is based on confidence level $\alpha$, allowing the joint-CP formulation using PRE to filter through predictions of the neural PDE solver. Upon being rejected, the initial conditions that led to those predictions could be provided to the expensive physics-based numerical PDE solver for further evaluation as indicated in Figure 1.

### v2

@@ Abstract
Simulating complex physical systems is crucial for understanding and predicting phenomena across diverse fields, such as fluid dynamics and heat transfer, as well as plasma physics and structural mechanics. Traditional approaches rely on solving partial differential equations (PDEs) using numerical methods, which are computationally expensive and often prohibitively slow for real-time applications or large-scale simulations. Neural PDEs have emerged as efficient alternatives to these costly numerical solvers, offering significant computational speed-ups. However, their lack of robust uncertainty quantification (UQ) limits deployment in critical applications. We introduce a model-agnostic, physics-informed conformal prediction (CP) framework that provides guaranteed uncertainty estimates without requiring labelled data. By utilising a physics-based approach, we can quantify and calibrate the model’s inconsistencies with the physics rather than the uncertainty arising from the data. Our approach utilises convolutional layers as finite-difference stencils and leverages physics residual errors as nonconformity scores, enabling data-free UQ with marginal and joint coverage guarantees across prediction domains for a range of complex PDEs. We further validate the efficacy of our method on neural PDE models for plasma modelling and shot design in fusion reactors.
@@ 1 Introduction
Numerical PDE solvers are essential tools in scientific and engineering simulations (Danabasoglu et al., 2020; Giudicelli et al., 2024), but their computational demands and environmental impact pose significant challenges (Horwitz, 2024). Machine learning approaches have emerged as efficient alternatives for modelling/emulating PDEs (Bertone et al., 2019; Karniadakis et al., 2021), successfully deployed across weather forecasting (Kochkov et al., 2024; Meyer et al., 2022; Giles et al., 2024), fluid dynamics (Jiang et al., 2020; Pfaff et al., 2021), and nuclear fusion applications (Poels et al., 2023; Carey et al., 2024; Gopakumar & Samaddar, 2020). Neural PDE solvers provide rapid approximations, but present a critical cost-accuracy trade-off. While generating outputs consistently, their solutions may violate physical constraints or produce misleading results with high confidence (Gopakumar et al., 2023a). A typical neural-PDE framework (see Figure 1) trains surrogate models on numerical simulation data to predict the evolution of PDE under various conditions, with uncertainty quantification (UQ) methods reverting to numerical solvers when predictions fail coverage thresholds. However, current UQ methods lack statistical guarantees (Zou et al., 2024), require extensive simulation data (Gopakumar et al., 2024a), or require architectural modifications (Abdar et al., 2021).
Figure 1: Neural PDE framework: Neural PDE solvers use data from traditional numerical solvers to quickly approximate PDEs across various conditions (shown by black arrows). To ensure reliability, these models incorporate uncertainty quantification (UQ) methods. If the predicted error exceeds a coverage threshold $\epsilon$, the numerical solver is utilised, further adding to the training data; otherwise, predictions are used as output (shown by red arrows). This paper focuses on developing a new UQ method for assessing confidence in neural PDE models, emphasising the grey-shaded region of the framework.
To address these limitations, we propose a framework combining PDE residuals over neural PDEs with Conformal Prediction (CP) to provide uncertainty estimates that guarantee coverage. Our approach evaluates Physics Residual Errors (PRE) from neural PDE solver predictions and performs calibration using marginal and joint CP formulations. Our method provides statistically valid coverage within the residual space (Vovk et al., 2005), offering error bounds against the violation of physical conservation laws. The framework is model-agnostic (applicable to all types of surrogate models) and does not require additional data generated from expensive numerical simulations. It yields interpretable uncertainty bounds indicating the model’s physics-guided inconsistencies, addressing the (over)confidence issue of neural PDE solvers (Zou et al., 2024). Our contributions:
@@ 2 Related Work
Recently, CP, as a method of performing UQ, has been gaining popularity for usage with spatio-temporal data (Sun, 2022). Several works have explored the inductive CP framework for spatial and sequential data (Stankeviciute et al., 2021; Xu & Xie, 2021; Xu et al., 2023), including in the operator space (Ma et al., 2024). Gopakumar et al. (2024a) extends the marginal-CP framework to pre-trained as well as to fine-tuned surrogate models for physical system modelling across an infinite-dimensional setting. Alternatively, error bounds for PDE surrogates have been devised by Gray et al. (2025) using set propagation to project the singular value decomposition of the prediction error onto the prediction space.
@@
The usage of PDE residuals under the guise of Physics-Informed Machine Learning (PIML) (Karniadakis et al., 2021) was made popular as an optimisation strategy for Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019) and has found application in optimising neural operators (Li et al., 2024) and soft/hard enforcement of the physical constraints to deep learning models (Du et al., 2024; Chalapathi et al., 2024). However, they have rarely been used as a tool for providing UQ to surrogate models, and where they have found application, UQ remained uncalibrated (Zhu et al., 2019). The majority of the literature in UQ for neural PDE solvers has been looking at Bayesian methods, such as dropout, Bayesian neural networks, and Monte Carlo methods (Geneva & Zabaras, 2020; Zou et al., 2024; Psaros et al., 2023), which lack guarantees or are computationally expensive.
@@ 3.1 Neural PDE Solvers
Here, $X$ defines the spatial domain bounded by $\Omega$, $[0,T]$ the temporal domain, and $D_{X}$ and $D_{t}$ the composite operators of the associated spatial and temporal derivatives. The PDE is further defined by the boundary condition $g$ and initial condition $a$, which can be parameterised by $\lambda$. The set of solutions of field variables is expressed as $u\in\mathcal{U}$.
Neural PDE solvers as surrogate models learn the behaviour governed by Equation 1 using a parameterised neural network $\mathcal{NN}_{\theta}$. Starting from the initial conditions, the network is trained to solve the spatio-temporal evolution of the fields given by $\Omega\,\cup\,[0,T]$. Neural operators $\mathcal{NO}_{\theta}$ are a special class of neural networks that learn the operator mapping from the function space of the PDE initial conditions $a\in\mathcal{A}$ to the function space of solutions $u\in\mathcal{U}$. A neural operator for solving an initial-value problem can be expressed as
@@
A Fourier Neural Operator (FNO) is an autoregressive neural operator that learns the spatio-temporal evolution of PDE solutions by leveraging the Fourier transform as the kernel integrator (Li et al., 2021). The field evolution is learned using tunable weight matrices of the network, parameterised directly in the Fourier space of the PDE solutions.
Since CP and our extension of it provide a post-hoc measure of quantifying the uncertainty of a neural PDE, it remains agnostic to model choice and training conditions. Considering the model independence of our approach, we restrict our experiments to modelling PDEs with an FNO. The FNO is chosen due to its cost-accuracy trade-off and efficiency as demonstrated by de Hoop et al. (2022) and Gopakumar et al. (2023b). CP over a range of neural-PDE solvers has been applied by Gopakumar et al. (2024a), who also demonstrate that the coverage guarantees are upheld irrespective of the model choice, further necessitating us to not experiment with various model architectures.
@@ 3.2 Conformal Prediction
Conventionally, nonconformity scores act on the model predictions and a labelled dataset (Kato et al., 2023). For deterministic models, they are often formulated as the Absolute Error Residual (AER) of the model predictions $\hat{f}(X)$ and targets $Y$. For probabilistic models, the score function (STD) is the absolute value of the z-score (with the prediction mean, standard deviation and target given as $\hat{f_{\mu}}(X),\hat{f_{\sigma}}(X),Y$ respectively. Having obtained a distribution of nonconformity scores $\hat{s}$ of the calibration dataset $(X_{i},Y_{i})_{i=1}^{n}$, a quantile $\hat{q}$ corresponding to the desired coverage $1-\alpha$ is estimated from its cumulative distribution function $F_{\hat{s}}$ (Papadopoulos, 2008):
@@
The quantile estimates the error bar associated with the desired coverage and is combined with the new prediction to obtain the prediction sets. The nonconformity score functions and their prediction sets for AER and STD are given in Table 1. Both AER and STD are data-intensive CP methods, requiring calibration data that converges to a beta distribution (Angelopoulos & Bates, 2023). This data dependence restricts their application to domains where sufficient data exists a priori or can be easily obtained
@@ 4 Physics Residual Error (PRE)
We introduce a novel data-free nonconformity score based directly on the PDE for surrogate models. The Physics Residual Error (PRE) is defined as the PDE residual (Saad & Schultz, 1986) estimated over the discretised PDE solution obtained from the surrogate model. For an abstract PDE as in Equation 1, the PDE residual is the evaluation of the composite differential operator $D$ over a field(s) $u$ under the influence of an external force $b$
$\mathcal{D}(u)-b=0,$ (7)
The PDE residual is treated as a score function by taking its L1 norm as indicated in Table 1. While well-defined PDEs have solutions obeying Equations 1, 2 and 3, numerical solutions often fail to converge to the true solution (Pinder, 2018). Neural PDEs, trained on approximate numerical data, are further prone to non-convergent predictions. In numerical analysis, the norm of the PDE residual is often used as a criterion for stability, convergence, and accuracy (Iserles, 2009). The PRE typically represents the violation of conservation laws associated with the physical system. Using the residual error as a nonconformity score quantifies the neural PDE solver’s non-convergence to the physical ground truth of the PDE. By further using conformal prediction, we obtain coverage bounds over this conservative, residual space with guaranteed coverage.
@@
The norm $\left|D(\mathcal{NO}_{\theta}(u))\right|$ of the residual operator itself provides a measure of UQ for the neural PDE. However, it is limited by the accuracy of the gradient estimation method and can become computationally expensive when exploring a vast solution space (Tolsma & Barton, 1998). By using the residual norm as a nonconformity score, we further calibrate the approximate physics residual error that an inexpensive and coarse differential operator obtains. CP using PRE provides statistically valid and guaranteed error bars across the PDE’s residual space, incorporating physical information into the calibration procedure and providing a calibrated measure of the physical misalignment of the surrogate model.
PRE as a nonconformity score enables data-free conformal prediction. The estimated scores rely only on the neural PDE predictions and not on the target as in AER and STD. The only criterion is that the calibration and prediction domains arise from exchangeable PDE conditions. As shown in Table 1, PRE gives “prediction sets” independent of the prediction inputs (see Appendix B for formalism). Traditional CP methods rely on calibration using labelled data to construct prediction intervals that contain the true values with a specified confidence level $\alpha$. These methods guarantee that the true solution will lie within the estimated error bounds based on this empirical calibration. Contrastingly, our CP-PRE formulation takes a fundamentally different approach. Instead of requiring target data for calibration, we leverage the unique property of PDEs where the true solution in the residual space exists at zero. This eliminates the need for empirical calibration data. Our method focuses on ensuring that predictions themselves fall within coverage bounds $\mathbb{C}^{\alpha}$, rather than guaranteeing that the true solution lies within these bounds. This allows us to validate prediction sets without access to ground truth data—a significant advantage over traditional CP approaches. We formalise this novel property in our theoretical framework presented in Appendix A.
@@ 4.1 Marginal-CP
The CP formulation was initially conceptualised for calibrating univariate functions with single-point outputs (Vovk et al., 2005). It has recently been extended to spatio-temporal data, with multi-dimensional outputs with an immutable tensor structure (Gopakumar et al., 2024a). Within such spatio-temporal settings, CP has been implemented to provide marginal coverage, i.e. the calibration procedure provides independent error bars for each cell within the spatio-temporal domain. For an output tensor $Y\in\mathbb{R}^{N_{x}\times N_{y}\times N_{t}}$, where $N_{x},N_{y},N_{t}$ represent the spatio-temporal discretisation of the domain, marginal-CP uses the non-conformity scores outlined in Tables 1, 1 and 1 across each cell of $Y$ to obtain error bars, which will be compliant with Equation 5 for each cell. Marginal-CP using PRE helps indicate spatio-temporal regions within individual predictions that lie outside the calibrated bounds of physics violation and require specific attention, treating those prediction regions with caution.
@@ 4.2 Joint-CP
The joint-CP formulation constructs a calibration procedure that provides coverage bands for multivariate functions. These coverage bands expand across the entire simulation domain $\Omega\times[0,T]$ (discretised as $\mathbb{R}^{N_{x}\times N_{y}\times N_{t}}$) rather than an individual cell within it. For a coverage band $\mathbb{C}^{\alpha}$, the joint-CP formulation ensures that $1-\alpha$ predictions/solutions lie within the bounds. For performing joint-CP, the non-conformity scores are modified to reflect the supremum of the score functions $\hat{s}$ in Tables 1, 1 and 1. They are modulated by the standard deviation $\sigma$ of the calibration scores (Diquigiovanni et al., 2021) to obtain prediction bands with varying widths based on local behaviour (Diquigiovanni et al., 2022). The modifications of the score functions and prediction sets to perform CP are given by
$\mathcal{S}_{\text{joint}}$ $=\sup_{X\in\Omega,\,t\in[0,T]}\bigg{(}\frac{\hat{s}}{\sigma(\hat{s})}\bigg{)},$ (8)
$\mathbb{C}^{\alpha}_{\text{joint}}$ $=p(X_{n+1})\pm\;\hat{q}^{\alpha}\cdot\sigma(\hat{s}),$ (9)
where $\hat{s}$ and $p(X_{n+1})$ are the formulations of the nonconformity scores and the prediction at $X_{n+1}$ used for marginal-CP as shown in Tables 1, 1 and 1. Joint-CP becomes particularly useful in identifying predictions that fail to fall within coverage, allowing us to accept or reject a prediction based on a predetermined probability. Similar to that demonstrated by Casella et al. (2004), our framework can perform acceptance-rejection using a CP-based criterion. The acceptance probability is based on confidence level $\alpha$. If a prediction is rejected, the PDE parameters that led to those predictions could be provided to the expensive physics-based numerical PDE solver for further evaluation as indicated in Figure 1.